Question # 1
Production point Pages typed Web pages created
A 0 4
B 40 3
C 70 2
D 90 1
E 100 0

(a) Jean can either type her term paper or create Web pages during the limited time she has available. The table above shows her PPF.
(i) Can Jean type 90 pages and create 2 Web pages?
(ii) Use the above numbers to calculate the opportunity cost of a typed page as she increases her time typing and decreases time creating a Web page.
1
Expert's answer
2020-05-27T10:03:15-0400

(i) The combination at work point D is beyond the scope of its PPF, so Jin cannot get 90 pages and create 2 web pages

(ii) Opportunity cost = (decrease in number of web pages) / (increase in number of type pages)

A to B = 4-3 and 40-0 = 1/40

B to C = 3-2 and 70-40 = 1/30

C to D = 2-1 and 90-70 = 1/20

D to E = 1-0 and 100-90 = 1/10